OpenStudy (anonymous):

Can you separate the equation below (y^2+2xy)y' + y^2 = 7x into the form of exact equation: M(x,y) dx + N(x,y) dy = 0 ? It seem like impossible ?

OpenStudy (anonymous):

as you should know: y'=dy/dx, so (y^2+2xy)(dy/dx)= 7x- y^2 then: (y^2+2xy)dy= (7x- y^2)dx

OpenStudy (anonymous):

Sorry, how did you get 7x ?

OpenStudy (anonymous):

but.... are we supposed to move the x to the right and y to the left?

OpenStudy (anonymous):

i mean there is still variable y in the dy equation.

OpenStudy (anonymous):

no: the finale shude be: (y^2+2xy)dy- (7x- y^2)dx =0

OpenStudy (anonymous):

ok... my bad....

OpenStudy (anonymous):

that's not important. You what exact equation is?

OpenStudy (anonymous):

I was confused with the separable equation.. hahaha

OpenStudy (anonymous):

ok, :)

OpenStudy (anonymous):

thanks a lot....

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I cannot give two best answer

OpenStudy (anonymous):

Thanks myko

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."